OpenStudy (anonymous):

write the standard form of the equation of the line describe. through:(2,4), perp. to y=-2/7x-5

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

y-4 = 7/2 (x-2) slope perp is negative reciprocal y = 7/2 x - 3